Please help me answer this question


A rectangular solid is constructed by putting together two cubes each of edge x ft. If the number
of cubic feet in the total volume is equal to the number of square feet in the total external surface area of this solid, what is the value of x?

Answer by edjones(8007)

surface area of each cube = 6x^2
5x^2*2=10x^2 suface area when joined.
.
x*x*2x=2x^3 volume when joined.
.
2x^3=10x^2
x^3=5x^2
x=5
.
Ed